the distance between singapore and tokyo is 3300 mi. on a certain wall map, this distance is represented be 11 in. the actual distance between mexico city and cairo is 7700 mi. how far apart are they on the same map?
Answer by Theo(13342)

same map same scale is assumed:
distance between singapore and tokyo is 3300 miles which is represented by 11 inches.
the ratio of miles to inches is equal to 3300 / 11 = 300.
the distance between mexico city and cairo is 7700 miles.
divide that by 300 and you get 7700 / 300 = 25.6666666666667 inches which is equivalent to 25 and 2/3 inches.
if you look at the ratio of the distances between the cities, you will see that the ratio is the same, regardless if your dealing with actual miles or with inches.
example:
7700/3300 = 2.333333333333 which is equivalent to 2 and 1/3.
25.6666666666666667 / 11 = 2.3333333333333333 which is equivalent to 2 and 1/3.
the assumption is that the miles to inches ratio remains the same regardless of where on the same map you are coming from and going to.